the earth's radius is about 6.4x10^6m. An object that has the mass of 20 kg is taken to height of 1.6x10^2 above the ground ? ( G = 6.67 x 10^-11 NM^2/kg^2 , ME = 6x10^24 kg ) numerical findout the mass of an object at this height and also how much does the object weights at this height ?
1
Expert's answer
2019-11-06T09:17:32-0500

The mass of an object at this height:


"m=20\ kg"

The weight:


"W=\frac{GmM_E}{(R_E+h)^2}"

"W=\frac{(6.67\cdot10^{-11})(20)(6\cdot10^{24})}{(6.4\cdot10^{6}+160)^2}=195.4\ N"